# From LLMs to LLM-based Agents for Software Engineering: A Survey of Current, Challenges and Future

### v1

@@ VI-B LLM-based Agents Tasks
Multi-agent collaboration and dialogue frameworks also demonstrated strong capabilities in both decision making and autonomous learning. [96] explores whether multi-agent discussions can enhance the reasoning abilities of LLMs. The proposed CMD framework simulates human group discussion processes, showing that multi-agent discussions can improve performance in commonsense knowledge and mathematical reasoning tasks without task-specific examples. Additionally, the study found that multi-agent discussions also correct common errors in single agents, such as judgment errors and the propagation of incorrect answers, thereby enhancing overall reasoning accuracy.  [97] researchers explored the potential of multi-modal large language models (MLLMs) like GPT4-Vision in enhancing agents’ autonomous decision-making processes. The paper introduce the PCA-EVAL benchmark, and evaluated multi-modal decision-making capabilities in areas such as autonomous driving, home assistants, and gaming. The results showed that GPT4-Vision exhibiting outstanding performance across the dimensions of perception, cognition, and action.
[98] proposes the Reflexion framework, a novel approach that strengthens learning through language feedback rather than traditional weight updates to avoid expensive re-train costs. The framework uses self-reflection and language feedback to help language agents learn from mistakes, significantly improving performance in decision-making, reasoning, and programming tasks. The Reflexion’s first-pass success rate on the HumanEval Python programming task increased from 80.1% to 91.0%, success rates in the ALFWorld decision-making task improved by 22%, and performance in the HotPotQA reasoning task increased by 14%. These results indicate that the Reflexion framework demonstrate the state-of-art performance in various tasks through self-reflection and language feedback.
@@
[99] proposes the AGENTVERSE multi-agent framework, designed to improve task completion efficiency and effectiveness through collaboration. The framework draws on human group dynamics by designing a collaborative system of expert agents that exhibit outstanding performance in tasks such as text understanding, reasoning, coding, and tool usage. Experiments showed that the AGENTVERSE framework performed well not only in independent task completion but also significantly improved performance through group collaboration especially in coding tasks where the framework use GPT-4 to be the brain of the agent groups. The framework also observed emergent behaviors in agents during collaboration, such as voluntary actions, conformity, and destructive behaviors, providing valuable insights for understanding and optimizing multi-agent systems.
@@
Integrating LLMs into agents allows for more complex task processing. [104] proposes a lightweight user-friendly library named AgentLite whic designed to simplify the development, prototyping and evaluation of task-oriented LLM-based agents systems. The main goal of the study is to enhance the capabilities and flexibility of LLM-based agents in various applications by introducing a flexible framework. This framework enhances task-solving capabilities through task decomposition and multi-agent coordination, using a hierarchical multi-agent coordination approach where a managing agent supervises the task execution of each agent. [105] introduces a framework, GPTSwarm, that represents LLM-based agents as computational graphs to unify existing prompt engineering techniques and introduces methods for optimizing these graphs to enhance agent performance. The study verifies the effectiveness of the framework through various benchmarks such as MMLU, Mini Crosswords, and HumanEval. The framework demonstrated significant performance improvements on the GAIA benchmark with an improvement margin of up to 90.2% compared to the best existing methods. Additionally, agents have shown strong capabilities in autonomous learning and decision-making in software engineering and security, which will be introduced in the subsequent Software Security section [106] [107] [108] [109].
@@ VI-C Analysis
Figure 6: Expel[35] Framework with Reflexion[98] in Experience Gathering
@@ VI-D Benchmarks
In the field of autonomous learning and decision-making, the benchmark datasets used by LLMs and LLM-based agents are quite similarly in the task handling and application requirements. We can gain a deeper understanding of the strengths and weaknesses of both approaches in different tasks and their application contexts. The specific dataset references, please see Table VII.
@@ VI-E Evaluation Metrics
TABLE VII: Evaluation Metrics in Autonomous Learning and Decision Making
@@ VII Software Design and Evaluation
LLMs in software design and evaluation also intersect extensively with autonomous learning and decision making, these two topics are interrelated fields. Software design needs to consider system adaptability and learning capabilities to handle dynamic environments, therefore design evaluations involving autonomous learning and decision making naturally become a focal point of intersection for these two topics. Many LLM techniques and methods find similar applications in both fields, for example LLMs based on reinforcement learning can be used for automated design decisions and evaluations, as well as for self-learning and optimization. Common applications of LLMs in software engineering involve fine-tuning models with prompt engineering techniques to continuously enhance performance particularly in software design and evaluation, more sample learning is often required to ensure that the model outputs align with user expectations [93] [102] [44] [111] [105] [96]. Additionally, requirement elicitation and specification in requirement engineering can also be considered part of software design and evaluation [51] [112]. This section reviews the main research achievements of LLMs in software design and evaluation in recent years, discussing their application scenarios and practical effects.
@@ VII-A LLMs Tasks
In recent years, there has been extensive research on the use of LLMs in tasks such as automation, optimization, and code understanding. ChatGPT has been widely utilized for various software engineering tasks and demonstrated excellent performance in tasks like log summarization, pronoun resolution, and code summarization, achieving a 100% success rate in both log summarization and pronoun resolution tasks [113]. However, its performance on tasks such as code review and vulnerability detection is relatively poor, which shows that it needs further improvement for more complex tasks. Another framework EvaluLLM addresses the limitations of traditional reference-based evaluation metrics (such as BLEU and ROUGE) by using LLMs to assess the quality of natural language generation (NLG) outputs [114]. The EvaluLLM introduces a new evaluation method that compares generative outputs in pairs and uses win rate metrics to measure model performance, this approach can simplifies the evaluation process also ensures consistency with human assessments, showcasing the broad application prospects of LLMs in generative tasks. Similarly, in the LLMs evaluation domain, LLM-based NLG Evaluation provides a review and classification of current LLMs used for NLG evaluation, the paper summarizes four main evaluation methods: LLM-derived metrics, prompt-based LLMs, fine-tuned LLMs, and human-LLM collaborative evaluations [115]. These methods demonstrate the potential of LLMs in evaluating generative outputs which also mention challenges such as the need for improved evaluation metrics and further exploration of human-LLM collaboration.
@@
There are also many novel application design with the LLMs which applied in the engineering design, one study explores strategies for software/hardware co-design to optimize LLMs and applies these strategies to design verification [116]. Through quantization, pruning, and operation-level optimization, this research demonstrates applications in high-level synthesis (HLS) design functionality verification, GPT-4 was used to generate high-level synthesis (HLS) designs containing predefined errors to create a dataset called Chrysalis, this dataset provides a valuable resource for evaluating and optimizing LLM-based HLS debugging assistants. The optimized LLM significantly improves inference performance, providing new possibilities for error detection and correction in the electronic design automation (EDA) field. In [117], the researchers introduces RaWi, a data-driven GUI prototyping approach. The framework allows users to retrieve GUIs from this repository, edit them, and create new high-fidelity prototypes quickly. The experiment conducted by comparing RaWi with a traditional GUI prototyping tool (Mockplus) to measure how quickly and effectively users can create prototypes. The result demonstrated that RaWi outperformed on multiple benchmarks, with 40% improvement on precision@k metric. This study proves the possibility of LLMs to improve the efficiency during prototyping phase of software design, which allows designers to quickly iterate on GUI designs, facilitating early detection of design flaws. With the new possibility brought by the LLMs, there has been much discussion in the education field, with researchers exploring the implications of the prevalence of large language models for education [118]. Study indicates that ChatGPT shows significant potential but some limitations in answering questions from software testing courses [119]. ChatGPT was able to answer about 77.5% of the questions and provided correct or partially correct answers 55.6% of the time. However, the correctness of its explanations was only 53.0%, indicating the need for further improvement in educational applications.
@@ VII-B LLM-based Agents Tasks
The application of LLM-based agents in software design and evaluation enhance the development efficiency and code quality, as well as showcase the broad applicability and immense potential of LLM-based agents in practical software engineering tasks. [120] explores the current capabilities, challenges, and opportunities of autonomous agents in software engineering. Study evaluate Auto-GPT’s performance across different stages of the software development lifecycle (SDLC), including software design, testing, and integration with GitHub, the paper finds that detailed contextual prompts significantly enhance agent performance in complex software engineering tasks which mentions the importance of context-rich prompts in reducing errors and improving efficiency, underscoring the potential of LLM-based agents to automate and optimize various SDLC tasks, thereby enhancing development efficiency. This paper also evaluate the limitation of the Auto-GPT, includes task or goal skipping, generating unnecessary code or files (hallucinations), repetitive or looping responses, lack of task completion verification mechanisms. These limitations can lead to incomplete workflows, inaccurate outputs, and unstable performance in practical applications.
@@
[121] introduces ChatDev, the first virtual chat-driven software development company, a concept of using LLMs not just for specific tasks but as central coordinators in a chat-based, multi-agent framework. this approach allows for more structured, efficient, and collaborative software development processes, exploring how chat-driven multi-agent systems can achieve efficient software design and evaluation, reduce code vulnerabilities, and enhance development efficiency and quality. Experiments show that ChatDev can design and generate software in an average of 409.84 seconds at a cost of only \$0.2967 while significantly reducing code vulnerabilities. This indicates that chat-based multi-agent frameworks capable to improve software development efficiency and quality. Another similar collaboration framework introduced by Microsoft research team,  [122] demonstrates the effectiveness of using LLMs, particularly ChatGPT as agent’s controllers to manage and execute various AI tasks. The HuggingGPT system that uses ChatGPT to orchestrate the execution of tasks by various AI models available in Hugging Face, the purpose is to test how effectively the system can handle complex AI tasks, including language, vision, and speech tasks, by executing appropriate models based on user requests. The innovation lies in using LLMs not just as tools for direct task execution but as central orchestrators that leverage existing AI models to fulfill complex tasks, This approach expands the practical applicability of LLMs beyond typical language tasks.  [123] proposes the LLMARENA benchmark framework to evaluate LLMs’ capabilities in dynamic multi-agent environments, the idea is similar to the ChatDev but innovates by shifting the focus from single-agent static tasks to dynamic and interactive multi-agent environments, providing a more realistic and challenging setting to assess the practical utility of LLMs, this approach mirrors real-world conditions where multiple agents (either AI or human) interact and collaborate. Experiments show that this framework can test LLMs’ spatial designing, strategic planning, and teamwork abilities in gaming environments, offering new possibilities and tools for designing and evaluating LLMs in multi-agent systems.
@@
[124] introduces the ”Flows” conceptual framework for structuring interactions between AI models and humans to improve reasoning and collaboration capabilities. The study present the idea of conceptualizing processes as independent, goal-driven entities that interact through standardized message-based interfaces, enabling a modular and extensible design. This approach is inherently concurrency-friendly and supports the development of complex nested AI interactions without having to manage complex dependencies. Experiments in competitive coding tasks show that the ”Flows” framework increases the AI model’s problem-solving rate by 21 percentage points and the human-AI collaboration rate by 54 percentage points. This demonstrates how modular design can enhance AI and human collaboration, thereby improving the software design and evaluation process.
[125] presents a new taxonomy to structurally understand and analyze LLM-integrated applications, providing new theories and methods for software design and evaluation. This taxonomy helps in understanding the integration of LLM components in software systems, laying a theoretical foundation for developing more effective and efficient LLM-integrated applications. Similarly, [126] explores the application of LLM-based agents in software maintenance tasks, improving code quality and reliability through a collaborative framework. This study should origin be categorized under the software maintenance domain but exhibit the iterative manner of the design structure. The framework utilize the task decomposition and multi-agent strategies to tackle complex engineering tasks that traditional one-shot methods cannot handle effectively, multiple agents can learn from each other, leading to improved software maintenance outcomes. Experiments show that multi-agent systems outperform single-agent systems in complex debugging tasks, indicating that this new framework can be applied in software design to provide safer architectures.

### v2

@@ VI-B LLM-based Agents Tasks
Multi-Agent Discussion and Reasoning Enhancement. Multi-agent collaboration and dialogue frameworks also demonstrated strong capabilities in both decision making and autonomous learning. [126] explores whether multi-agent discussions can enhance the reasoning abilities of LLMs. The proposed CMD framework simulates human group discussion processes, showing that multi-agent discussions can improve performance in commonsense knowledge and mathematical reasoning tasks without task-specific examples. Additionally, the study found that multi-agent discussions also correct common errors in single agents, such as judgment errors and the propagation of incorrect answers, thereby enhancing overall reasoning accuracy.  [127] researchers explored the potential of multi-modal large language models (MLLMs) like GPT4-Vision in enhancing agents’ autonomous decision-making processes. The paper introduce the PCA-EVAL benchmark, and evaluated multi-modal decision-making capabilities in areas such as autonomous driving, home assistants, and gaming. The results showed that GPT4-Vision exhibiting outstanding performance across the dimensions of perception, cognition, and action.
Self-Reflection and Experience-Based Learning. [59] proposes the Reflexion framework, a novel approach that strengthens learning through language feedback rather than traditional weight updates to avoid expensive re-train costs. The framework uses self-reflection and language feedback to help language agents learn from mistakes, significantly improving performance in decision-making, reasoning, and programming tasks. The Reflexion’s first-pass success rate on the HumanEval Python programming task increased from 80.1% to 91.0%, success rates in the ALFWorld decision-making task improved by 22%, and performance in the HotPotQA reasoning task increased by 14%. These results indicate that the Reflexion framework demonstrate the state-of-art performance in various tasks through self-reflection and language feedback.
@@
Multi-Agent Collaboration and Emergent Behavior. [58] proposes the AGENTVERSE multi-agent framework, designed to improve task completion efficiency and effectiveness through collaboration. The framework draws on human group dynamics by designing a collaborative system of expert agents that exhibit outstanding performance in tasks such as text understanding, reasoning, coding, and tool usage. Experiments showed that the AGENTVERSE framework performed well not only in independent task completion but also significantly improved performance through group collaboration especially in coding tasks where the framework use GPT-4 to be the brain of the agent groups. The framework also observed emergent behaviors in agents during collaboration, such as voluntary actions, conformity, and destructive behaviors, providing valuable insights for understanding and optimizing multi-agent systems.
@@
Tooling and Engineering Frameworks for Agents. Integrating LLMs into agents allows for more complex task processing. [132] proposes a lightweight user-friendly library named AgentLite whic designed to simplify the development, prototyping and evaluation of task-oriented LLM-based agents systems. The main goal of the study is to enhance the capabilities and flexibility of LLM-based agents in various applications by introducing a flexible framework. This framework enhances task-solving capabilities through task decomposition and multi-agent coordination, using a hierarchical multi-agent coordination approach where a managing agent supervises the task execution of each agent. [133] introduces a framework, GPTSwarm, that represents LLM-based agents as computational graphs to unify existing prompt engineering techniques and introduces methods for optimizing these graphs to enhance agent performance. The study verifies the effectiveness of the framework through various benchmarks such as MMLU, Mini Crosswords, and HumanEval. The framework demonstrated significant performance improvements on the GAIA benchmark with an improvement margin of up to 90.2% compared to the best existing methods.  [134] presents Iris, an LLM-based virtual tutor embedded within the Artemis learning platform, designed to support students in programming exercises. The system integrates context-awareness by automatically gathering code, problem descriptions, and test feedback, enabling it to provide relevant and personalized assistance without requiring manual input from students, empirical results show that Iris effectively enhances learning engagement and comprehension, while students maintain confidence to work independently in exams. This study represents a practical instantiation of LLM-based tool frameworks in real-world software education environments.  [56] presents GoEx, an open-source runtime specifically designed for executing actions generated by LLM agents in a secure, reversible, and isolated manner. GoEx targets the infrastructure-level trust and reliability challenges of autonomous LLM deployment in real-world systems. The core concept proposed is post-facto validation, allowing agents to act first while enabling undo and damage-confinement mechanisms for safety. GoEx supports actions across RESTful APIs, databases, and file systems, providing containerized execution environments, credential vaults, reversion sets, and blast-radius control. Additionally, agents have shown strong capabilities in autonomous learning and decision-making in software engineering and security, which will be introduced in the subsequent Software Security section VI  [135, 136, 137, 138].
Bridging Code Generation and Autonomous Learning. While the majority of works in this section emphasize decision-making, feedback, or learning strategies, several frameworks also demonstrate strong capabilities in real-world code generation tasks, thus forming natural intersections between Autonomous Learning and Decision Making and Code Generation and Software Development. For example, CodeAct [114], this approach goes beyond traditional code generation by enabling LLMs to reason through control flows and conditionals, showing strong decision-making ability. Similarly, CODEAGENT [54] and the multi-agent Agile framework [110] both automates not only code creation but also iterative task decomposition and coordination, blending collaborative decision-making with procedural learning. Another representative cross-topic work is ToolCoder [139], which reformulates tool learning as a code generation problem and systematically integrates software engineering principles into the agent planning and execution pipeline. Rather than relying on natural language chains-of-thought, ToolCoder converts user tasks into structured Python function scaffolds, enabling modular task decomposition, tool selection, and precise action planning. Its modular code-centric design demonstrates strong capabilities in both software-oriented code execution and adaptive decision making under uncertainty, bridging the gap between traditional tool usage and agent-style autonomous behavior.
@@ VI-C Analysis
Figure 7: Expel[40] Framework with Reflexion[59] in Experience Gathering.
@@ VI-D Benchmarks
In the field of autonomous learning and decision-making, the benchmark datasets used by LLMs and LLM-based agents are quite similar in task handling and application requirements. We can gain a deeper understanding of the strengths and weaknesses of both approaches in different tasks and their application contexts. The specific dataset references, please see Table VIII.
@@ VI-E Evaluation Metrics
TABLE VIII: Evaluation Metrics in Autonomous Learning and Decision Making
@@
 [56] None N/A (Design Perspective and Architectural Analysis) Yes
 [134] Artemis Programming Exercises Dataset Correctness, Engagement,
Learning Confidence Yes
 [139] RestBench (TMDB, Spotify), API-Bank, ToolAlpaca Success%, Path%, Accuracy%,
Correctness%, Procedure, Response Yes
@@ VII Software Design and Evaluation
LLMs in software design and evaluation also intersect extensively with autonomous learning and decision making, these two topics are interrelated fields. Software design needs to consider system adaptability and learning capabilities to handle dynamic environments, therefore design evaluations involving autonomous learning and decision making naturally become a focal point of intersection for these two topics. Many LLM techniques and methods find similar applications in both fields, for example LLMs based on reinforcement learning can be used for automated design decisions and evaluations, as well as for self-learning and optimization. Common applications of LLMs in software engineering involve fine-tuning models with prompt engineering techniques to continuously enhance performance particularly in software design and evaluation, more sample learning is often required to ensure that the model outputs align with user expectations [123, 130, 49, 141, 133, 126]. Additionally, requirement elicitation and specification in requirement engineering can also be considered part of software design and evaluation [64, 142]. This section reviews the main research achievements of LLMs in software design and evaluation in recent years, discussing their application scenarios and practical effects.
@@ VII-A LLMs Tasks
Software Task Evaluation and Assessment. In recent years, there has been extensive research on the use of LLMs in tasks such as automation, optimization, and code understanding. ChatGPT has been widely utilized for various software engineering tasks and demonstrated excellent performance in tasks like log summarization, pronoun resolution, and code summarization, achieving a 100% success rate in both log summarization and pronoun resolution tasks [143]. However, its performance on tasks such as code review and vulnerability detection is relatively poor, which shows that it needs further improvement for more complex tasks. Another framework EvaluLLM addresses the limitations of traditional reference-based evaluation metrics (such as BLEU and ROUGE) by using LLMs to assess the quality of natural language generation (NLG) outputs [26]. The EvaluLLM introduces a new evaluation method that compares generative outputs in pairs and uses win rate metrics to measure model performance, this approach can simplifies the evaluation process also ensures consistency with human assessments, showcasing the broad application prospects of LLMs in generative tasks. Similarly, in the LLMs evaluation domain, LLM-based NLG Evaluation provides a review and classification of current LLMs used for NLG evaluation, the paper summarizes four main evaluation methods: LLM-derived metrics, prompt-based LLMs, fine-tuned LLMs, and human-LLM collaborative evaluations [144]. These methods demonstrate the potential of LLMs in evaluating generative outputs which also mention challenges such as the need for improved evaluation metrics and further exploration of human-LLM collaboration.
@@
Software and Hardware Design Automation. There are also many novel application design with the LLMs which applied in the engineering design, one study explores strategies for software/hardware co-design to optimize LLMs and applies these strategies to design verification [145]. Through quantization, pruning, and operation-level optimization, this research demonstrates applications in high-level synthesis (HLS) design functionality verification, GPT-4 was used to generate high-level synthesis (HLS) designs containing predefined errors to create a dataset called Chrysalis, this dataset provides a valuable resource for evaluating and optimizing LLM-based HLS debugging assistants. The optimized LLM significantly improves inference performance, providing new possibilities for error detection and correction in the electronic design automation (EDA) field. In [146], the researchers introduces RaWi, a data-driven GUI prototyping approach. The framework allows users to retrieve GUIs from this repository, edit them, and create new high-fidelity prototypes quickly. The experiment conducted by comparing RaWi with a traditional GUI prototyping tool (Mockplus) to measure how quickly and effectively users can create prototypes. The result demonstrated that RaWi outperformed on multiple benchmarks, with 40% improvement on precision@k metric. This study proves the possibility of LLMs to improve the efficiency during prototyping phase of software design, which allows designers to quickly iterate on GUI designs, facilitating early detection of design flaws.
Software Engineering Education. With the new possibility brought by the LLMs, there has been much discussion in the education field, with researchers exploring the implications of the prevalence of large language models for education [147]. Study indicates that ChatGPT shows significant potential but some limitations in answering questions from software testing courses [148]. ChatGPT was able to answer about 77.5% of the questions and provided correct or partially correct answers 55.6% of the time. However, the correctness of its explanations was only 53.0%, indicating the need for further improvement in educational applications.
@@ VII-B LLM-based Agents Tasks
Assessment and Limitation Studies in LLM-Based Agents. The application of LLM-based agents in software design and evaluation enhance the development efficiency and code quality, as well as showcase the broad applicability and immense potential of LLM-based agents in practical software engineering tasks. [149] explores the current capabilities, challenges, and opportunities of autonomous agents in software engineering. Study evaluate Auto-GPT’s performance across different stages of the software development lifecycle (SDLC), including software design, testing, and integration with GitHub, the paper finds that detailed contextual prompts significantly enhance agent performance in complex software engineering tasks which mentions the importance of context-rich prompts in reducing errors and improving efficiency, underscoring the potential of LLM-based agents to automate and optimize various SDLC tasks, thereby enhancing development efficiency. This paper also evaluate the limitation of the Auto-GPT, includes task or goal skipping, generating unnecessary code or files (hallucinations), repetitive or looping responses, lack of task completion verification mechanisms. These limitations can lead to incomplete workflows, inaccurate outputs, and unstable performance in practical applications.
@@
Multi-Agent Collaboration for Software Engineering.  [109] introduces ChatDev, the first virtual chat-driven software development company, a concept of using LLMs not just for specific tasks but as central coordinators in a chat-based, multi-agent framework. this approach allows for more structured, efficient, and collaborative software development processes, exploring how chat-driven multi-agent systems can achieve efficient software design and evaluation, reduce code vulnerabilities, and enhance development efficiency and quality. Experiments show that ChatDev can design and generate software in an average of 409.84 seconds at a cost of only \$0.2967 while significantly reducing code vulnerabilities. This indicates that chat-based multi-agent frameworks capable to improve software development efficiency and quality.  [150] proposes the LLMARENA benchmark framework to evaluate LLMs’ capabilities in dynamic multi-agent environments, the idea is similar to the ChatDev but innovates by shifting the focus from single-agent static tasks to dynamic and interactive multi-agent environments, providing a more realistic and challenging setting to assess the practical utility of LLMs, this approach mirrors real-world conditions where multiple agents (either AI or human) interact and collaborate. Experiments show that this framework can test LLMs’ spatial designing, strategic planning, and teamwork abilities in gaming environments, offering new possibilities and tools for designing and evaluating LLMs in multi-agent systems. Similarly, [151] explores the application of LLM-based agents in software maintenance tasks, improving code quality and reliability through a collaborative framework. This study should origin be categorized under the software maintenance domain but exhibit the iterative manner of the design structure. The framework utilize the task decomposition and multi-agent strategies to tackle complex engineering tasks that traditional one-shot methods cannot handle effectively, multiple agents can learn from each other, leading to improved software maintenance outcomes. Experiments show that multi-agent systems outperform single-agent systems in complex debugging tasks, indicating that this new framework can be applied in software design to provide safer architectures.
@@
Agent-Orchestrated Software Design and Development.  [152] introduces the ”Flows” conceptual framework for structuring interactions between AI models and humans to improve reasoning and collaboration capabilities. The study present the idea of conceptualizing processes as independent, goal-driven entities that interact through standardized message-based interfaces, enabling a modular and extensible design. This approach is inherently concurrency-friendly and supports the development of complex nested AI interactions without having to manage complex dependencies. Experiments in competitive coding tasks show that the ”Flows” framework increases the AI model’s problem-solving rate by 21 percentage points and the human-AI collaboration rate by 54 percentage points. This demonstrates how modular design can enhance AI and human collaboration, thereby improving the software design and evaluation process. Another similar collaboration framework introduced by Microsoft research team,  [153] demonstrates the effectiveness of using LLMs, particularly ChatGPT as agent’s controllers to manage and execute various AI tasks. The HuggingGPT system that uses ChatGPT to orchestrate the execution of tasks by various AI models available in Hugging Face, the purpose is to test how effectively the system can handle complex AI tasks, including language, vision, and speech tasks, by executing appropriate models based on user requests. The innovation lies in using LLMs not just as tools for direct task execution but as central orchestrators that leverage existing AI models to fulfill complex tasks, This approach expands the practical applicability of LLMs beyond typical language tasks.  [154] presents a new taxonomy to structurally understand and analyze LLM-integrated applications, providing new theories and methods for software design and evaluation. This taxonomy helps in understanding the integration of LLM components in software systems, laying a theoretical foundation for developing more effective and efficient LLM-integrated applications.